# SincVAE: a New Approach to Improve Anomaly Detection on EEG Data Using SincNet and Variational Autoencoder

### v1

@@ Abstract
Over the past few decades, electroencephalography (EEG) monitoring has become a pivotal tool for diagnosing neurological disorders, particularly for detecting seizures. Epilepsy, one of the most prevalent neurological diseases worldwide, affects approximately the 1 % of the population. These patients face significant risks, underscoring the need for reliable, continuous seizure monitoring in daily life. Most of the techniques discussed in the literature rely on supervised Machine Learning (ML) methods. However, the challenge of accurately labeling variations in epileptic EEG waveforms complicates the use of these approaches. Additionally, the rarity of ictal events introduces an high imbalancing within the data, which could lead to poor prediction performance in supervised learning approaches. Instead, a semi-supervised approach allows to train the model only on data not containing seizures, thus avoiding the issues related to the data imbalancing. This work proposes a semi-supervised approach for detecting epileptic seizures from EEG data, utilizing a novel Deep Learning-based method called SincVAE. This proposal incorporates the learning of an ad-hoc array of bandpass filter as a first layer of a Variational Autoencoder (VAE), potentially eliminating the preprocessing stage where informative band frequencies are identified and isolated. Results indicate that SincVAE improves seizure detection in EEG data and is capable of identifying early seizures during the preictal stage as well as monitoring patients throughout the postictal stage.
Keywords:
Variational Autoencoders SincNet Anomaly Detection Seizure Detection Brain Computer Interfaces1 Introduction
@@
In the last years, the information technology growth had a significant impact on our daily life. In particular, its impact is evident by observing the quantity of data produced every day. These vast datasets provide a snapshot of the entities under observation, offering valuable insights for companies and organizations. These insights not only enhance understanding but also furnish competitive advantages. Consequently, it is crucial that these datasets are meticolously processed [1]. Anomalies (known also as outliers, deviants or rare events in some context [2]) represent unique behaviors within observed phenomena that can significantly influence the data generation process [2, 3]. The presence of anomalies in data during analysis can be dangerous, as they may lead to erroneous conclusions during data interpretation. Consequently, it is crucial for the analysts that these anomalies are meticulously identified and properly addressed of both before and during the analysis process. The increasing interest in identifying and analyzing anomalies led scientists to isolate this problem into the active and dedicated research field of anomaly detection.
@@
In most instances, phenomena are monitored over time using time series data. Within this framework, not every outlier is pertinent to an analysis. For instance, some outliers could be attributed to sensor transmission errors or other sources of noise, while others might represent unusual phenomena, such as those observed in fraud detection scenarios [2]. In the former scenario, outliers can be eliminated or corrected to enhance data quality. Conversely, in the latter case, these outliers transform into anomalies of interest that needs special attention as they often provide significant and crucial insights across diverse application domains [2]. For instance, anomalies in credit card transaction records may indicate potential fraud or identity theft [4]. Similarly, an unusual pattern in network traffic could suggest that a compromised computer is transmitting confidential information to an unauthorized destination [5]. Additionally, anomalies detected in sensor data from civil infrastructure might signal structural damage [6, 7].
@@
Significantly, healthcare represents a domain where anomaly detection can have a profound impact [8, 9, 10, 11, 12]. For example, the ability to detect abnormal physiological data through these techniques can expedite emergency responses and provide new insights into the progression of medical conditions, greatly influencing everyday life. In [13], the authors presented a methodology for identifying anomalies in heart rate data, leveraging on its value as a noninvasive indicator of health concerns and physical activity. Meanwhile, the study in [14] explores the potential of anomaly detection within healthcare analytics, specifically through IoT systems.
@@
Numerous studies have explored detecting epileptic seizures as abnormal brain activity using electroencephalography (EEG) data, applicable in various settings, including Brain-Computer Interfaces (BCI) [15]. In such contexts, accurately identifying seizures through EEG can trigger hardware interventions designed to assist patients and enhance their quality of life [16, 17, 18]. Epilepsy is a chronic central nervous system disorder characterized by recurrent seizures, affecting approximately 1 % of the global population [19, 20]. Seizures manifest as temporary disruptions in brain electrical activity, leading to symptoms like attention lapses, memory gaps, sensory hallucinations, or full-body convulsions. Despite treatment with various anti-epileptic drugs, about one-third of affected individuals frequently experience seizures that are challenging to control. These seizures significantly increase the risk of injury, limit personal independence and mobility, and can result in social and economic difficulties [21, 22]. The brain activity of individuals with epilepsy can be categorized into four states: regular brain activity (interictal), brain activity preceding the seizure (preictal), brain activity during the seizure (ictal), and brain activity immediately following a seizure (postictal) [23].
@@
Among various methodologies [24, 25], Machine Learning (ML) techniques provide systematic approaches for extracting insights from the vast amounts of data. These techniques enable researchers to explore the anomaly detection landscape, proposing solutions to diverse scenarios. In particular, Deep Learning (DL) methods have consistently outperformed traditional Machine Learning approaches in the last decades. This success is largely due to their ability to extract intricate patterns within complex, high-dimensional datasets [26, 27]. As a result, DL has taken a leading role in various fields that leverage data-driven strategies, especially in the development of anomaly detection methodologies [28, 29, 30]. A significant portion of DL approaches for anomaly detection relies on Autoencoder (AE) architectures [31, 32, 33]. AEs are neural networks comprising two main components: an encoder, which compresses the input into a latent representation, and a decoder, which reconstructs that representation back into the input space. In anomaly detection, a well-established strategy involves training Autoencoders (AEs) specifically to minimize reconstruction errors for normal data instances. This technique results in higher reconstruction errors when processing anomalous data, making these errors a useful anomaly score. When combined with a user-defined decision rule, this score becomes an effective tool for classifying data as normal or anomalous [6]. AEs architectures can be realized in various forms, including Variational Autoencoders (VAE), that is a generative model where the encoder and the decoder does not represent a functional mapping as in standard AEs [34]. Due to their promising performances, there is a growing interest in using generative models to identify anomalies [35, 36, 8, 37] that, within the context of AEs, has been manifested through the use of VAEs [38, 39, 40, 41].
@@
VAEs can be implemented using a variety of processing strategies documented in literature, including Multilayer Perceptrons (MLP) [42] and Recurrent Neural Networks (RNN) [43]. Convolutional Neural Networks (CNNs), widely utilized for processing time series data [44, 45, 46, 47, 48], operate through a series of trainable filters. These filters are inspired by the biological mechanisms of visual perception, enabling the recognition of informative patterns from the input data [49]. In the field of speaker recognition, Ravanelli and Bengio introduced SincNet [50], a CNN whose filters are structured as a learnable array of parametrized sinc functions that are designed to operate as bandpass filters.
@@
Numerous ML methods have been applied to detect epileptic seizures, with the aim of not only mitigate risks for patients but also enhance their ability to seek timely assistance and reduce the likelihood of injury [23, 51, 52, 53]. Although the physiological activity involved in epilepsy is inherently multiclass, many studies, such as [21, 22], approach seizure detection as a binary supervised classification problem. In this framework, the two classes to be identified are seizure activity (ictal) and non-seizure activity (interictal). This reduction to two classes is due to the challenges and impracticalities associated with an expert’s ability to identify and label transitional states between ictal and interictal states. Conversely, having an expert categorize brain electrical activity into seizure and non-seizure aligns with standard clinical protocols [22].
This work proposes SincVAE, a DL architecture that introduces SincNet in the VAE framework to process EEG data for seizure detection in a semi-supervised setting. Bandpass filters are crucial for isolating meaningful frequency bands from input signals, especially in BCI applications [54, 55, 56, 57]. However, the application of bandpass filters typically involves two stages: (i) selecting well-known informative band frequencies pertinent to the specific context, and (ii) conducting an analysis where the analyst identifies and isolates relevant frequency bands from the data. SincNet offers an efficient solution in this process, providing a compact and precise method to develop custom bandpass filters optimized for specific applications. Applications of seizure detection could benefit from this improvement since it could allow to enhance (or eventually, eliminate), preliminary phases dedicated to the study and extraction of frequency bands from the data and, thus, achieve a more refined and precise extraction of frequency bands, leading to an acceleration and improvement of the whole processing pipeline development. A semi-supervised approach is particularly advantageous as it allows to fit a model on training sets that contain only normal instances [2, 58, 59, 6]. This choice is motivated by the flexibility this approach offers, especially in scenarios with imbalanced datasets, where anomalous data points are scarce or even absent. This is particularly relevant in healthcare, where the prevalence of normal data contrasts with the rarity of abnormal data.
@@ 2 Related Works
Khan et al. in [60] shows a novel seizure detection method by integrating AEs with traditional classifiers in an hybrid model. Specifically, the AE is used to extract features from the input data through its encoder. These latent representations are then fed into a classifier, such as an Support Vector Machine (SVM) [61] or k-Nearest Neighbors (k-NN) [61], to perform a supervised classification. Yuan et al. in [62] proposed a novel approach that employs an AE model to extract multi-view features from multi-channel EEG data. Then, the features extracted are used for supervised seizure detection. Abdelhameed et al. in [63] proposed a methodology based on convolutional VAE to extract features from EEG input data with the goal of eliminating the need of an engineered feature extraction phase previous to the model fitting. Then, the extracted features are fed in a supervised classifier to detect seizures. The same authors in [64] improved the feature extraction phase by using a two-dimensional Deep Convolutional Autoencoder (2D-DCAE). The extracted features are then used to train a neural network-based classifier for seizure detection in a supervised manner. In [65] instead, the same authors proposed a methodology based on a convolutional VAE trained in a supervised manner to perform simultaneously automatic feature learning and classification on the data latent representations. Daoud et al. in [66] compared two methodologies, both having the automatic features extraction as main goal. The first method utilizes a deep convolutional AE, where features extracted from the encoder are classified using a Multilayer Perceptron (MLP) classifier. The second method consists in an unsupervised pipeline integrating a deep convolutional VAE with the K-Means [61] clustering algorithm fitted on the latent representations of data. Wang et al. in [67] introduced the Residual Convolution VAE (RCVAE) method to extract features from EEG recordings. The extracted features are used to train a supervised neural network classifier for the seizure detection. The same research group in [68] proposed an improved version called Residual Convolution VAE with Randomly Translation Strategy (RTS-RCVAE) to solve issues related to introduce data augmentation strategies. Wen et al. in [69] proposed the AE-CDNN method to extract the feature previous to a supervised classification stage performed using MLPs. Similarly, Shoeibi et al. in [70] employed autoencoders for dimensionality reduction. Latent representations of data were used to fit several methods to classify seizures including adaptive neuro-fuzzy inference system (ANFIS) and its variants.
@@
Most of the works on seizure detection leverages on AEs or VAEs primarily for feature extraction, which are then utilized in conjunction with supervised methods for classifying EEG data. There exists a relatively small subset of methodologies that rely exclusively on AEs and VAEs for the entire process of seizure detection, using only the reconstruction error metrics for the classification stage. Huang et al. in [71] used AEs for feature extraction in epilepsy detection, comparing their performance against traditional Principal Component Analysis (PCA) [61]. Then, the authors employed three metrics, i.e. original-to-reconstructed signal ratio (ORSR), MSE and Cosine Similarity (CS), to evaluate the signal reconstruction and identify these metrics as sensitive indicators for epilepsy. Also, the authors utilize permutation importance and SHapley Additive exPlanations (SHAP) [72] for model interpretability, confirming the better efficacy and rationale of the AE-based feature extraction compared to the PCA one. The authors in [73] leveraged on a three-layered convolutional VAEs trained exclusively on non-seizure EEG recordings to detect seizures. Reconstruction error was used to identify seizure activity. In particular, they adopted the median reconstruction error as a metric to distinguish between seizure and non-seizure events. You et al. in [12] used a VAE to model the latent representations of non-seizure EEG signals. They then used deviations from these baseline representations in conjunction with reconstruction loss to devise a personalized anomaly score for each patient. De Sousa et al. in [74] used AEs and VAEs to detect Interictal Epileptiform Discharges (IEDs) by treating these events as anomalies within EEG data. The comparative analysis in their study revealed that VAEs outperformed traditional AEs, likely due to their enhanced ability to model the distribution of EEG data and handle anomalies more effectively. Potter et al. in [75] proposed an architecture based on AE with a transformer encoder to reconstruct EEG recordings of non-seizure activity. Then, reconstruction error is served as anomaly score to detect EEG recordings containing seizures,

### v2

@@ Abstract
Over the past few decades, electroencephalography monitoring has become a pivotal tool for diagnosing neurological disorders, particularly for detecting seizures. Epilepsy, one of the most prevalent neurological diseases worldwide, affects approximately of the population. These patients face significant risks, underscoring the need for reliable, continuous seizure monitoring in daily life. Most of the techniques discussed in the literature rely on supervised machine learning methods. However, the challenge of accurately labeling variations in epileptic electroencephalography waveforms complicates the use of these approaches. Additionally, the rarity of ictal events introduces a high imbalance within the data, which could lead to poor prediction performance in supervised learning approaches. Instead, a semi-supervised approach allows training the model only on data that does not contain seizures, thus avoiding the issues related to the data imbalance. This work introduces a semi-supervised approach for detecting epileptic seizures from electroencephalography data based on a novel deep learning-based method called SincVAE. This method integrates SincNet, designed to learn an ad-hoc array of bandpass filters, as the first layer of a variational autoencoder, potentially eliminating the preprocessing stage where informative frequency bands are identified and isolated. Experimental evaluations on the Bonn and CHB-MIT datasets indicate that SincVAE improves seizure detection in electroencephalography data, with the capability to identify early seizures during the preictal stage and monitor patients throughout the postictal stage.
1 Introduction
@@
In recent years, the growth of information technology has had a significant impact on our daily life. In particular, its impact is evident by observing the quantity of data produced every day. These vast datasets provide a snapshot of the entities under observation, offering valuable insights for companies and organizations. These insights not only enhance understanding but also furnish competitive advantages. Consequently, it is crucial that these datasets are meticulously processed [1]. Anomalies (also known as outliers, deviants or rare events in some context [2]) represent unique behaviors within observed phenomena that can significantly influence the data generation process [2, 3]. The presence of anomalies in data during analysis can be dangerous, as they may lead to erroneous conclusions during data interpretation. Consequently, it is crucial for analysts to meticulously identify and properly address these anomalies both before and during the analysis process. The increasing interest in identifying and analyzing anomalies led scientists to isolate this problem into the active and dedicated research field of anomaly detection.
@@
In most instances, phenomena are monitored over time using time series data. Within this framework, not every outlier is pertinent to an analysis. For instance, some outliers could be attributed to sensor transmission errors or other sources of noise, while others might represent unusual phenomena, such as those observed in fraud detection scenarios [2]. In the former case, outliers can be eliminated or corrected to enhance data quality. Conversely, in the latter case, these outliers become anomalies of interest, often providing significant and crucial insights across diverse application domains [2]. For instance, anomalies in credit card transaction records may indicate potential fraud or identity theft [4]. Similarly, an unusual pattern in network traffic could suggest that a compromised computer is transmitting confidential information to an unauthorized destination [5]. Additionally, anomalies detected in sensor data from civil infrastructure might signal structural damages [6, 7].
@@
Significantly, healthcare represents a domain where anomaly detection can have a profound impact [8, 9, 10, 11, 12]. For example, the ability to detect abnormal physiological data through these techniques can expedite emergency responses and provide new insights into the progression of medical conditions, greatly influencing everyday life. In [13], the authors presented a methodology for identifying anomalies in heart rate data, leveraging its value as a noninvasive indicator of health concerns and physical activity. Meanwhile, the study in [14] explores the potential of anomaly detection within healthcare analytics, specifically through IoT systems.
@@
Numerous studies have explored detecting epileptic seizures as abnormal brain activity using electroencephalography (EEG) data, which are applicable in various settings, including brain-computer interfaces (BCI) [15, 16]. In such contexts, accurately identifying seizures through EEG can trigger hardware interventions designed to assist patients and improve their quality of life [17, 18, 19]. Epilepsy is a chronic central nervous system disorder characterized by recurrent seizures, affecting approximately of the global population [20, 21]. Seizures manifest as temporary disruptions in brain electrical activity, leading to symptoms like attention lapses, memory gaps, sensory hallucinations, or full-body convulsions. Despite treatment with various anti-epileptic drugs, about one-third of affected individuals frequently experience seizures that are challenging to control. These seizures significantly increase the risk of injury, limit personal independence and mobility, and can result in social and economic difficulties [22, 23]. The brain activity of individuals with epilepsy can be categorized into four states: regular brain activity (interictal), brain activity preceding the seizure (preictal), brain activity during the seizure (ictal), and brain activity immediately following a seizure (postictal) [24].
@@
Among various methodologies [25, 26], machine learning (ML) techniques provide systematic approaches for extracting insights from a vast amount of data. These techniques enable researchers to explore the anomaly detection landscape, and propose solutions to diverse scenarios. In particular, deep learning (DL) methods have consistently outperformed traditional ML approaches in the last decades. This success is largely due to their ability to extract intricate patterns within complex, high-dimensional datasets [27, 28]. As a result, DL has played a leading role in various fields that leverage data-driven strategies, especially in the development of anomaly detection methodologies [29, 30, 31]. A significant portion of DL approaches for anomaly detection relies on autoencoder (AE) architectures [32, 33, 34]. AEs are neural networks comprising two main components: an encoder, which compresses the input into a latent representation, and a decoder, which reconstructs that representation back to the input space. In anomaly detection, a well-established strategy involves training AEs specifically to minimize reconstruction errors for normal data instances, effectively framing it as a semi-supervised problem [6]. This technique results in higher reconstruction errors when processing anomalous data, making these errors a useful anomaly score. When combined with a user-defined decision rule, this score becomes an effective tool for classifying data as normal or anomalous [6]. Using this methodology, AEs effectively handle variations in normal data while being less affected by the specific characteristics of rare and diverse anomalies. Moreover, their semi-supervised approach eliminates the need for labeled anomaly data during training, making them well-suited for real-world scenarios where such labels are frequently limited or unavailable. AEs architectures can be realized in various forms, including variational autoencoder (VAE), which is a generative model where the encoder and the decoder do not represent a functional mapping as in standard AEs [35]. Due to their promising performance, there is a growing interest in using generative models to identify anomalies [36, 37, 8, 38] that, within the context of AEs, has been manifested through the use of VAEs [39, 40, 41, 42].
@@
VAEs can be implemented using a variety of processing strategies documented in the literature, including multilayer perceptrons (MLP) [43] and recurrent neural networks (RNN) [44]. Convolutional neural networks (CNNs), widely utilized for processing time series data [45, 46, 47, 48, 49], operate through a series of trainable filters. These filters are inspired by the biological mechanisms of visual perception, enabling the recognition of informative patterns from the input data [50]. In the field of speaker recognition, Ravanelli and Bengio introduced SincNet [51], a CNN whose filters are structured as a learnable array of parametrized sinc functions that are designed to operate as bandpass filters.
@@
Numerous ML methods have been applied to detect epileptic seizures, with the aim of not only mitigating risks for patients but also enhancing their ability to seek timely assistance and reduce the likelihood of injury [24, 52, 53, 54]. Although the physiological activity involved in epilepsy is inherently multi-class, according to the four brain activity states introduced above, many studies, such as [22, 23], approach seizure detection as a binary supervised classification problem, where the two classes to be identified are seizure activity (ictal) and non-seizure activity (interictal). This reduction to two classes is due to the challenges and impracticalities associated with an expert’s ability to identify and label transitional states between ictal and interictal states, whereas an expert can categorize brain electrical activity as seizure or non-seizure, adhering to current standard clinical protocols [23].
Furthermore, epilepsy detection could significantly benefit from a semi-supervised learning approach, especially given the scarcity of seizure data in contrast to the abundance of non-seizure data. For a patient known to suffer from epilepsy, data collected during non-seizure states can be easily obtained and can provide a substantial foundation for building a ML model capable of accurately detecting seizures when they occur.
This work proposes SincVAE, a DL architecture that introduces SincNet in the VAE framework to process EEG data for seizure detection in a semi-supervised setting.
Bandpass filters are crucial for isolating meaningful frequency bands from input signals, especially in BCI applications [55, 56, 57, 58].
However, the application of bandpass filters typically involves two stages: (i) selecting well-known informative band frequencies pertinent to the specific context, and (ii) conducting an analysis where the analyst identifies and isolates relevant frequency bands from the data.
SincNet offers an efficient method in this process, providing a compact and precise solution to develop custom bandpass filters optimized for specific applications through the training stage.
Applications of seizure detection could benefit from this improvement since it could allow to enhance (or eventually, eliminate), preliminary phases dedicated to the study and extraction of frequency bands from the data and, thus, achieve a more refined and precise extraction of frequency bands, leading to an acceleration and improvement of the whole processing pipeline development.
Furthermore, SincNet filters act as bandpass filters, and after the training stage, they exhibit inherent interpretability [51], offering analysts valuable insights into the frequency bands relevant to the seizure detection on the training dataset. The challenge of interpreting DL models is widely acknowledged in the literature, as their over-parameterized, black-box nature often obscures the rationale behind their predictions [59]. Consequently, numerous algorithms have been proposed to enhance the interpretability of DL models by analyzing their components and understanding their behavior during inference.
In particular, various strategies have been proposed to address these issues, including post-hoc explanation methods (e.g., saliency maps, SHAP, LIME), model simplification, and the design of inherently interpretable architectures. In most conventional deep learning models, the learned configurations, such as convolutional filters or dense layer weights, lack clear structure and are not directly interpretable, making it difficult to understand what the model has actually learned [60].
In this context, SincNet allows to have a notable advantage, as its filters design inherently provides interpretability without requiring the application of additional algorithms. This built-in transparency enables a more direct understanding of the model’s processing, offering valuable insights into the specific frequency bands involved, which is particularly important for clinical applications.
Additionally, the signal decomposition performed by SincNet enhances the VAE’s capabilities in detecting and extracting meaningful features that are representative of the non-anomalous condition during the training phase. This improvement allows for more accurate representation of the non-anomalous state, thereby emphasizing the distinction between non-anomalous and anomalous patterns in the data.
A semi-supervised approach [61, 62, 63, 6, 7] obtained through the application of a VAE network is particularly advantageous as it allows to fit a model on training sets that contain only non-anomalous instances [2, 64, 65, 6], in this context represented by interictal brain activity.
This choice is motivated by the flexibility this approach offers, especially in scenarios with imbalanced datasets, where anomalous data points are scarce or even absent.
Indeed, data imbalancing is a well-known challenge in healthcare problems, and various proposals in the literature have been made to address it [66, 67, 68].
This is particularly relevant in healthcare, where the prevalence of non-anomalous data contrasts with the rarity of abnormal data.
This paper is organized as follows: Section 2 provides an overview of existing works on AEs and VAEs for seizure detection from EEG data; Section 3 introduces the proposed method, beginning with an overview of the SincNet and VAE frameworks; Section 4 details the experimental assessment; Section 5 presents the results of the experimental evaluation; and Section 6 concludes the paper, discussing potential future directions.
@@ 2 Related Works
Khan et al. in [69] show a novel seizure detection method by integrating AEs with traditional classifiers in an hybrid model. Specifically, the AE is used to extract features from the input data through its encoder. These latent representations are then fed into a classifier, such as a support vector machine (SVM) [70] or k-nearest neighbors (k-NN) [70], to perform a supervised classification. Yuan et al. in [71] proposed a novel approach that employs an AE model to extract multi-view features from multi-channel EEG data. Then, the extracted features are then used for supervised seizure detection. Abdelhameed et al. in [72] proposed a methodology based on convolutional VAE to extract features from EEG input data with the goal of eliminating the need for an engineered feature extraction phase previous to the model fitting. Then, the extracted features are fed in a supervised classifier to detect seizures. The same authors in [73] improved the feature extraction phase by using a two-dimensional deep convolutional AE (2D-DCAE). The extracted features are then used to train a neural network-based classifier for seizure detection in a supervised manner. In [74] instead, the same authors proposed a methodology based on a convolutional VAE trained in a supervised manner to perform simultaneously automatic feature learning and classification on the data latent representations. Daoud et al. in [75] compared two methodologies, where the main goal was automatic feature extraction. The first method utilizes a deep convolutional AE, where features extracted from the encoder are classified using a MLP classifier. The second method consists in an unsupervised pipeline integrating a deep convolutional VAE with the k-means [70] clustering algorithm fitted on the latent representations of data. Wang et al. in [76] introduced the residual convolution VAE (RCVAE) method to extract features from EEG recordings. The extracted features are used to train a supervised neural network classifier for seizure detection. The same research group in [77] proposed an improved version called residual convolution VAE with randomly translation strategy (RTS-RCVAE) to solve issues related to the introduction of data augmentation strategies. Wen et al. in [78] proposed the AE-CDNN method to extract the features prior to a supervised classification stage performed using MLPs. Similarly, Shoeibi et al. in [79] employed AEs for dimensionality reduction. Latent representations of data were used to fit several methods to classify seizures including the adaptive neuro-fuzzy inference system (ANFIS) and its variants. Huang et al. in [80] used AEs for feature extraction in epilepsy detection, comparing their performance against traditional principal component analysis (PCA) [70]. Then, the authors employed three metrics, namely original-to-reconstructed signal ratio (ORSR), mean squared error (MSE) and cosine similarity (CS), to evaluate the signal reconstruction and identify these metrics as sensitive indicators for epilepsy. Also, the authors utilize permutation importance and Shapley additive explanations (SHAP) [81] for model interpretability, confirming the better efficacy and rationale of the AE-based feature extraction compared to the PCA one.
@@
Most of the works on seizure detection rely on AEs or VAEs primarily for feature extraction, which are then utilized in conjunction with supervised methods for classifying EEG data. There exists a relatively small subset of methodologies that rely exclusively on AEs and VAEs for the entire process of seizure detection, using only the reconstruction error metrics for the classification stage. The authors in [82] relied on a three-layered convolutional VAEs trained exclusively on non-seizure EEG recordings to detect seizures. Reconstruction error was used to identify seizure activity. In particular, they adopted the median reconstruction error as a metric to distinguish between seizure and non-seizure events. You et al. in [12] used a VAE to model the latent representations of non-seizure EEG signals. They then used deviations from these baseline representations, along with reconstruction loss, to devise a personalized anomaly score for each patient. De Sousa et al. in [83] used AEs and VAEs to detect interictal epileptiform discharges (IEDs) by treating these events as anomalies within EEG data. The comparative analysis in their study revealed that VAEs outperformed traditional AEs, likely due to their enhanced ability to model the distribution of EEG data and handle anomalies more effectively. Potter et al. in [84] proposed an architecture based on AE with a transformer encoder to reconstruct EEG recordings of non-seizure activity. Then, reconstruction error served as anomaly score to detect EEG recordings containing seizures.